OpenStudy (anonymous):

If the APR of a savings account is 4.8% and interest is compounded monthly, what is the approximate APY of the account? A. 10.49% B. 4.91% C. 1.05% D. 4%

OpenStudy (anonymous):

Hey braycub, I believe I have your answer. This is only my second attempt at answering, so I may do something amiss, but here goes! APR = Annual Percentage Rate; whereas APY = Annual Percentage Yield. As such, the APY is always going to be a touch higher than APR unless the instrument (the savings account in this case) only compounds yearly. With just this knowledge, you can guess the answer to be B. 4.91% because APY and APR are always close, just that APY is a touch higher. To prove it, you use this formula, or a financial calculator: APY = [((1\times(1+APR \div n)^{n})-1\] n is equal to the number of times the instrument in question compounds in a year, so 12 in this case. Ie., APY = (1 * (1+0.048/12)^n) - 1 = 1+0.04907 - 1 = 0.0491, or as a percentage, 4.91%.

OpenStudy (anonymous):

Ack, I need to clean up my post, but I don't know how to edit yet >_< . Alas. It's correct, but could be cleaner looking.

OpenStudy (anonymous):

ALRIGHT THANKS BUDDY